2C4H10 + 13O2  →  10H2O + 8CO2

What is the molar mass of O2 ? 


Question 1 options:


416.00 g/mol


208.00 g/mol


16.00 mol/g


32.00 mol/

Given the balanced equation:

2C4H10 + 13O2  →  10H2O + 8CO2

What is the molar mass of C4H10 ? 


Question 2 options:


116.28 g/mol


58.14 g/mol


116.28 mol/

Question 3 (2 points)

 


Saved







Given the balanced equation:

2C4H10 + 13O2  →  10H2O + 8CO2

If you are given 150 grams of C4H10 ,How many moles would that be? 


Question 3 options:


5.16 mol


2.58 mol


0.19 mol


0.39 mo

Question 4 (3 points)

 








Given the balanced equation:

2C4H10 + 13O2  →  10H2O + 8CO2

If you are given 150 grams of C4H10 ,How many moles of O2 would you need to completely react with the C4H10 ? 


Question 4 options:


5.16 mol


8.39 mol


16.77 mol


0.06 mo

Question 5 (4 points)

 








Given the balanced equation:

2C4H10 + 13O2  →  10H2O + 8CO2

If you are given 150 grams of C4H10 ,How many grams of O2 would you need to completely react with the C4H10 ? 


Question 5 options:


150


975 grams


4800


537 gram


1
Expert's answer
2021-02-18T07:46:33-0500

Ans:- (1)

Molar mass of Oxygen is 16 g/mol

Mass of "O_2" = 16"\times" 2 "\Rightarrow" 32g/mol

So the molar mass of "O_2" according to the balanced equation is =13"\times" 32 "\Rightarrow"416.00g/mol

Ans:- (2)

Molecular mass of "C_4H_{10}" = 58.14g/mol

So the molar mass of "C_4H_{10}" according to balanced equation is = 58.14"\times" 2"\Rightarrow" 116.28 g/mol

Ans:- (3)

Moles of "C_4H_{10}" = "\dfrac{150}{58.14}" "\Rightarrow" 2.58 mol

So, the moles of "C_4H_{10}" according to stoichiometry is 2.58"\times" 2 "\Rightarrow" 5.16 mol

Ans:-(4)

According to Stoichiometry

Since, 2 moles of "C_4H_{10}" = 13 moles of "O_2" reacted

moles of "O_2" reacted completely to produce product= "\dfrac{13\times 2.58}{2}=16.77mol"

Ans:-(5)

Grams of "O_2" = moles of "O_2 \times" molar mass of "O_2"

"\Rightarrow""{16.77}\times32"

"\Rightarrow" 537 gm